OpenStudy (anonymous):

Prove that square root of 5 times 5^{3/2}=25

geerky42 (geerky42):

Try to use exponents product rules: \(\Huge a^m\cdot a^n=a^{m+n}\)

OpenStudy (aum):

\[\large \sqrt{5} \times 5^{3/2} = 5^{1/2} \times 5^{3/2} = 5^{3/2 + 1/2} = 5^{2} = 25 \]

OpenStudy (anonymous):

thanks guys

OpenStudy (aum):

yw.